OpenStudy (anonymous):

The temperature was 65 degrees at daybreak. Then it dropped 2 degrees per hour until dusk. If x is the number of hours since daybreak, this equation represents the temperature during the day: y = -2x + 65 OR y = 65 - 2x. Use the equation to predict the temperature 6 hours later.

OpenStudy (anonymous):

A. 53 B. 63 C. 61 D. 59

OpenStudy (anonymous):

Did you attempt to solve? If they want they temperature for 6 hours later, than 6 becomes your "X". y=-2(6)+65 y=-12+65 y=53

OpenStudy (anonymous):

Thank you so much!